# Modeling when and how physics PhD students search for a research group: the role of interests and prior research experiences in timely group integration

### v2

@@ IV.3 Students with narrower interests join research groups more quickly
Having established a spectrum of categories describing students’ degree of certainty regarding their research interests, this section examines differences in when students with varying levels of interest joined their research groups. To do this, we sorted students into groups based on the highest level of interest they expressed before starting graduate school ($N=15$ in Broad interests, $N=12$ in Interest in subfield and method, and $N=13$ in Knowing exact research). We then split each interest group by when they committed to a research group and compared student counts.
@@
As illustrated in Figure 4, we observe that students who expressed having broad research interests upon entering graduate school were more likely to join a group in their second semester of graduate school or later. Figure 4a) shows that 12 of 15 students who were categorized as having broad interests found their research groups during the second semester or later, compared to 3 of 15 who joined in their first semester or before. The green bar of 4a) shows that of the 22 students who joined a research group during or after their second semester, 12 of 22 (54%) had been coded as having broad research interests. This is compared to the red bar, which illustrates that 3 of 18 (17%) students who committed to a group first semester or before were coded as having broad interests. Comparing Figure 4a) to Figs. 4b) and 4c), we see that the time to join a research group decreases as students’ certainty of interest increases. Figure 4b) shows 33% ($N=6$) of students who joined first semester or before expressed interest in subfield and method, compared to 27% ($N=6$) who joined after first semester. Meanwhile, in Figure 4c), we see that students who came into graduate school with strong research interests tended to commit to a research group early. Indeed, 50% ($N=9$) of first semester or earlier joiners were students who described knowing the exact research topic they wanted to pursue. This is compared to just 18% ($N=4$) of the students who found a group during their second semester or later.
@@
In addition to sorting students by when they joined a research group, Figure 4 also indicates whether the student had switched out of a different research group prior to joining their current lab. Of the four students in Figure 4c) who came in with strong interests but joined later, three (Benjamin, Kwan, and Tabitha) had switched research groups. This shows how these students had indeed joined a research group quickly upon entering graduate school, but discovered that their groups did not fit them well. For example, Benjamin began working with a research group his senior summer in order to pursue his interest in fusion research, but quickly realized that he did not like his work or the group’s overall culture as much as he thought he would. He recalled, “I don’t know. I just wasn’t super excited about the project… and there was like, Zoom meetings with a lot of cranky old scientists that were like yelling matches. It was kind of a nasty environment.” Thus, neither the research topic or group environment fit Benjamin’s expectations. However, since Benjamin had this experience so early in his graduate career, he was able to switch groups during his 2nd semester into one that better suited his interests.
@@ IV.3.1 Access to undergraduate research clarifies research interests
Taken as a whole, these results demonstrate that degree of certainty with regard to research interests is an important element in an explanatory model characterizing how and why students join research groups. Moreover, they indicate some of the ways that a student’s background and access to advanced coursework and undergraduate research opportunities might impact their search for a research group down the road. Research on the impacts of undergraduate research experiences has shown that students tend to constrain their options for future work to topics with which they are familiar [83, 84]. Thus, opportunities to engage with modern research areas in physics as an undergraduate make paths into those areas seem more open, and simultaneously open doors to pursuing them in the future. Elena’s comment that working in high energy experiment had been “already been laid out” for her based on her previous research experience echoes these findings.
@@
In alignment with prior literature, students in our sample with stronger interests often cited their undergraduate research experiences as a major factor impacting their interest development. Desiree described her undergraduate research working in high energy theory as the “key event” in her interest development, saying “That was when I was like, okay, this seems like what I want to do.” Similarly, Fiona noted that she “did some research where I did my bachelor’s, and then I also did the two REUs. And from that I knew what I was interested in was mainly the theory side of stuff, and specifically stars.”
Despite the fact that undergraduate research experiences are often critical for students’ interest development, access to undergraduate research is lower for traditionally underrepresented groups in physics, particularly ethnic minorities and first-generation college students [85]. One first-generation student in our sample, Carmen, relayed how “In undergrad, I mean, I am a first-generation student. And I, as many people, struggled with my course load… I struggled to find someone who could advise me in research, and I applied to REU programs over the summer. My own department and externally. The first time I did that, I did not receive any offers.” The following year, Carmen earned a research position in their home department doing astrophysics, indicating that this experience was the reason they were able to pursue a graduate degree: “I probably wouldn’t be here to talking with you today [if not for the REU].”
@@
Still, Carmen felt as though “there were people who had, in my opinion, more resources and more exposure to the field at an early age than I did. And so I still didn’t know what I wanted to do.” Thus, Carmen entered graduate school unsure of the research they wanted to pursue and needing to find a group during their first year of graduate study. However, as detailed in our previous work [31], this opened the door to a number of new challenges as Carmen searched for a group while trying to adjust to their new environment and taking graduate-level courses. At points during this process, Carmen considered leaving the program as these challenges accrued. Hence, Carmen’s experience is a microcosm of the broader message of this section, that inequitable access to formative research experiences can have long-term effects on the first-year graduate experience. We will return to Carmen’s story in more detail in Section IV.5.
@@
Access to advanced coursework and electives is one factor that shapes students’ interests, helping them narrow their research interests sooner. Although Desiree’s interest in high energy theory was confirmed during her research experience, it initially stemmed from taking advanced electives: “During my junior and senior years, I started taking more electives… particle physics, cosmology, things that are not in like the main courseload… [topics] you literally never hear about in undergrad, unless you decide to go take a class on it.” Similarly, Matias cited the graduate-level condensed matter course he took as a senior as being a pivotal influence, as he strongly disliked the undergraduate version. He recalled, “That professor was teaching graduate solid state, and I was like oh shit, here we go again. But I enrolled. And then when we looked at it from a graduate perspective, I started to like it. I was like, wait, okay, actually this is kind of cool. We looked into topics in a deeper sense you know?” Yet students from smaller departments, particularly those without a graduate program, might not offer such courses at all.
Despite the predominant trend that students who express a high degree of certainty in their interests tend to join groups more quickly, we see also observe several deviations from this tendency. Four students who described entered graduate school knowing the exact research they wanted to pursue ended up joining a group sometime after their second semester. For one of these students, funding issues for her preferred advisor meant she had to wait until the summer of her first year to begin work with him. The other three began working in their preferred research right away, but negative experiences drove them to switch groups. These students’ narratives are discussed at length in Section IV.5, where we more deeply examine the difficulties and concerns that students had while searching for a group.
@@ IV.4 Early connections with prospective advisors help students join groups sooner
Another major factor that Lovitts cited as predicting success in the PhD was students’ proactivity in searching for a research group. Specifically, she found that in the period before arriving at graduate school, students who collected information beyond just the school level were more likely to persist. Students who contacted faculty directly or went for on-campus visits were more likely to complete a PhD than students who relied more on the school’s reputation or looking at faculty websites: “the closer the source of information to the most interpersonal interactional level (faculty), the larger and more significant the difference between completers and noncompleters becomes.” In this section, we illustrate the variation in when students start earnestly looking into research groups, and show how students who engage in the search process earlier tend to join a group more quickly.
@@
Students with broader interests than Wendy indicated that their criteria for narrowing down schools was not as stringent. Isaac said that while applying to schools, “I knew vaguely what area of physics I was interested in… I guess, at the time, particle physics, and maybe like quantum field theory, or something. So my thought was just, you know what, let’s see if there’s anyone out there doing that.” Online resources also played a big role at this time for Isaac, who said he would “go to [the physics department’s] page, look at the professors, and just click on people who seemed interesting.” A common refrain from students at this point in their process was that they wanted to identify programs with several groups whose research seemed to align with their interests. For instance, Blake said that she was drawn to her graduate program because “because there’s a lot of research in what I wanted to do here,” which was appealing because “I didn’t have a super specific focus I wanted within optics” (Senior year).
@@
Many students ($N=28$) had excerpts coded in the School level alignment category, and students were equally distributed over time to join a group (72% who joined 1st semester or before and 68% who joined after first semester). This is indicative of how regardless of whether students were sure of what research they wanted to pursue, they still went through this process. Still, this means 12 students did not discuss looking at their school’s research at all. While several students may have simply omitted this from their narratives, some students specifically noted that they did not filter schools based on research. For instance, two students attended the same graduate school where they went to undergrad, and never contemplated going elsewhere. Another student chose which graduate schools to apply for based on whether they offered a fee waiver, figuring they would sort out their research interests once they arrived. Other students primarily filtered graduate schools based on their location and proximity to family. Meanwhile, several students strongly coupled their choice of graduate school to an individual advisor, and therefore never talked about considering the research going on in the broader department.
@@
Indivdual group level alignment Excerpts with student attempts to determine research interest congruence with just one or a few groups were coded as Evaluating individual group level alignment. While the school level alignment phase commonly had students looking at school websites for help, this phase most commonly characterized by talking to prospective advisors and graduate students directly. In both stages, students reflected often on their undergraduate research experiences for help deciding whether a research topic seemed interesting. In sum, excerpts from $N=34$ unique students were coded under Evaluating individual group alignment, again indicating that students went through this process regardless of interest level.
@@
Students often described becoming more focused on particular groups once they arrived in graduate school. For instance, Valerie recalled beginning to “go between research group to research group, all these different meetings… definitely comparing them against each other, and then comparing them against the good internship experience I had.” Hassan said that it was “first semester when I slowly started getting more into, like the in-depth portion of the research that he did.” This involved meeting directly with his prospective advisor: “I went and met with him. We briefly talked about what he does, I got a tour of the lab.” Jiya emphasized that one of the primary benefits of being on campus with prospective advisors was that she could ask them directly about their research, saying “I want to hear it from them, like how they describe it.”
@@
Within the category of looking at individual groups, several students described specifically wanting to evaluate research at the Project level. Excerpts coded under this category show how students sought to understand more precisely what they would work on in the lab if they were to join, and what to expect in their day-to-day tasks. For example, after being accepted to his top choice graduate school in his Senior year of undergrad, Alex recalled asking professors, “Where do you see your research going in the next three to six years? Is the work that you’re doing now going to continue, or am I going to join your group and find that you’re pivoting 180 degrees? And you know, now you’re not doing the work that I joined your group to do” (Senior year). Matias similarly wanted to know what the “day-to-day” of his work would be like in a prospective group, which he was able to answer by earning a fellowship to work there the summer before graduate school. He said, “I would say if I didn’t have that opportunity, I would still be very uncertain right now.” Comparatively fewer students ($N=15$) made statements coded in under this sub-category, suggesting that it may not be a topic that students looking for groups consider as often or feel that they are able to evaluate before deciding to join a group.
@@
In total, $N=24$ students talked about evaluating their interest in individual groups before graduate school, while $N=11$ only talked about looking at research groups after arriving in graduate school. Figure 5 plots these two subsets of students, split by when they joined a research group. We see that students who joined a research group during or before their 1st semester were more likely to have evaluated individual groups or projects before they arrived in graduate school than students who found a group second semester or later. 72% ($N=13$) of students who committed to a research group in their first semester or earlier had started the process of evaluating individual research groups before graduate school, compared to 45% ($N=10$) of students who joined second semester or later. Only 11% ($N=2$) of the earlier joiners indicated that they had not already been looking at specific groups prior to arriving in graduate school. On the other hand, 41% ($N=9$) of those who joined second semester or later did not begin closely looking into individual groups until starting their programs. This illustrates the variation in when students start earnestly looking into research groups, and shows how students who engage in the search process earlier tend to join a group more quickly.
Figure 5 also highlights the subset of students in our sample who switched research groups, shown in dark green. The two students represented by the dark green bar on the left are Benjamin and Tabitha, both of whom were also categorized as knowing the exact research they wanted to pursue. Benjamin was guided toward his research group by his undergraduate advisor, who was a collaborator with his prospective graduate advisor; Tabitha’s prospective advisor contacted her directly to recruit her to his group. However, unlike many of the other students who found their groups through strong connections to their undergraduate research, both students found that their graduate labs did not fit them well and switched.
@@ IV.4.1 Connections from undergraduate experiences help students search earlier
The resources that students used while looking at schools or contacting individual research groups varied. Websites and prior experiences helped to figure out which schools might have interesting research, but students often had difficulty gleaning more specific details about a professor’s research solely from online resources. For that, other sources of help, such as one-on-one conversations with faculty, were significantly more valuable. At the same time, websites are significantly more accessible and less intimidating than meetings with graduate faculty.
@@
Prior to arriving in graduate school, setting up meetings with prospective advisors with whom students are not co-located is more difficult and requires more effort. Matias discussed the lack of time associated with finishing undergraduate coursework at the same time as applying to graduate school: “I barely had time because I’m finishing up my semester, I did a double major so every semester was like 20 units, so I was like, up to here with work. And so it was really, it’s really hard to set time aside to look into these things, you know? Also… I was like, I’m not even gonna get in. So what’s the point of looking into it?” Some discussed anxiety associated with reaching out to faculty. For example, Jiya indicated that she struggled with being able to “put myself out there” and start talking to faculty. Isaac recalled that “I guess the biggest question I had was, if I were to summarize, how to find someone I’m interested in and then how to talk to them and have them not think I’m an idiot.”
@@
Some students’ experiences put them in a better position to navigate these difficulties. For instance, Felipe’s undergraduate research advisor helped him make connections with his eventual group in graduate school: “One of [my mentors] encouraged me to apply here. Because he knew the people in the research group I was interested in working with, so it was like, seems like a nice fit.” Felipe leveraged his undergraduate advisor’s connections to break the ice with his eventual graduate advisor. Similarly, Alex said that emailing faculty was easy after his REU, which he described as “a program whose sole purpose was to help students get into [the school, so I] had those connections essentially made for me. So when I wanted to take advantage, all I needed to do was send an email, basically no introduction necessary.” The networking opportunities that undergraduate research provided allowed some students to more efficiently get in touch with individual groups. Meanwhile, Selena cited her background at a small school as a distinct disadvantage during her search. “My school was small enough, we didn’t have connections to these places I was applying. So it wasn’t like I could say to the person I’m working with in my school, can you reach out to someone there? Can you help me find someone? So it was very much like Google, and a lot of clicking, was how I tried to piece it together.”
@@
These results suggest some of the resources used to more closely evaluate research groups (e.g., meetings with faculty) may not be as easily accessible to some students as others. Some are less likely to have the time or connections to look beyond the school level while narrowing down groups, Or they simply may not know whether certain activities, such as reaching out to faculty before starting their graduate program, are appropriate.

### v3

@@ IV.3 Students with narrower interests join research groups more quickly
Having established a spectrum students’ research interest certainty, this section examines differences in when students with varying levels of interest joined their research groups. To do this, we sorted students into groups based on the highest level of interest they expressed before starting graduate school ($N=15$ in Broad interests, $N=12$ in Interest in subfield and method, and $N=13$ in Knowing exact research). We then split each interest group by when they committed to a research group and compared student counts.
@@
As illustrated in Figure 4, we observe that students in our sample who expressed having broad research interests upon entering graduate school were more likely to join a group in their second semester of graduate school or later. Figure 4a) shows that 12 of 15 students who were categorized as having broad interests found their research groups during the second semester or later, compared to 3 of 15 who joined in their first semester or before. The green bar of 4a) also demonstrates that of all 22 students who joined a research group during or after their second semester, 12 of them (54%) had been coded as having broad research interests. This is compared to the red bar, which illustrates that of all 18 students who committed to a group first semester or before, only 3 (17%) were coded as having broad interests.
Comparing Figure 4a) to Figs. 4b) and 4c), we see that the time to join a research group decreases as certainty of interest increases. Figure 4b) shows 33% ($N=6$) of students who joined first semester or before expressed interest in subfield and method, compared to 27% ($N=6$) who joined after first semester. Meanwhile, in Figure 4c), we see that students who came into graduate school with strong research interests tended to commit to a research group early. Indeed, 50% ($N=9$) of all first semester or earlier joiners were students who described knowing the exact research topic they wanted to pursue. This is compared to just 18% ($N=4$) of the students who found a group during their second semester or later.
@@
In addition to sorting students by when they joined a research group, Figure 4 also indicates whether the student had switched out of a different research group prior to joining their current lab. Of the four students in Figure 4c) who came in with strong interests but joined later, three (Benjamin, Kwan, and Tabitha) had switched research groups. This shows how these students had indeed joined a research group quickly upon entering graduate school, but discovered that their groups did not fit them well. For example, Benjamin began working with a research group his senior summer in order to pursue his interest in fusion research, but quickly realized that he did not like his work or the group’s overall culture as much as he thought he would. He recalled, “I don’t know. I just wasn’t super excited about the project… and there was like, Zoom meetings with a lot of cranky old scientists that were like yelling matches. It was kind of a nasty environment.” Thus, neither the research topic or group environment fit Benjamin’s expectations. However, since Benjamin had this experience so early in his graduate career, he was able to switch groups during his 2nd semester into one that better suited his interests. Tabitha and Kwan’s narratives are discussed at length in Section IV.5, where we more deeply examine the difficulties and concerns that students had while searching for a group.
@@ IV.3.1 Access to undergraduate research clarifies research interests
Aligning with prior literature discussed in Section I on the impact of undergraduate research, students in our sample with stronger interests often cited their undergraduate research experiences as a major factor impacting their interest development. For example, Elena’s comment that working in high energy experiment as an undergraduate made her feel that her path had “already been laid out.” Similarly, Desiree described her undergraduate research working in high energy theory as the “key event” in her interest development, saying “That was when I was like, okay, this seems like what I want to do.” Fiona noted that she “did some research where I did my bachelor’s, and then I also did the two REUs. And from that I knew what I was interested in was mainly the theory side of stuff, and specifically stars.” All of these quotes illustrate how opportunities to engage with modern research areas in physics as an undergraduate make paths into those areas seem more open, and simultaneously open doors to pursuing them in the future.
However, several students in our data cited struggles with finding a path into undergraduate research, which they associated with factors such as their demographic backgrounds or the small size of their undergraduate institution. These findings are also in alignment with prior literature regarding access to undergraduate research experiences that were discussed in Section I. For example, Hassan was interested in topics related to quantum computing, but noted that his small undergraduate institution did not offer enough research opportunities for him to determine the kind of research he wanted to do. He said, “I didn’t really have a lot of research background. So like, in undergrad, I got involved with like one project, which was small scale given the school that I attended. But I was not really sure what I wanted to do.”
@@
One first-generation student in our sample, Carmen, relayed how “In undergrad, I mean, I am a first-generation student. And I, as many people, struggled with my course load… I struggled to find someone who could advise me in research, and I applied to REU programs over the summer. My own department and externally. The first time I did that, I did not receive any offers.” The following year, Carmen earned a research position in their home department doing astrophysics, indicating that this experience was the reason they were able to pursue a graduate degree: “I probably wouldn’t be here to talking with you today [if not for the REU].” Still, Carmen felt as though “there were people who had, in my opinion, more resources and more exposure to the field at an early age than I did. And so I still didn’t know what I wanted to do.” Carmen entered graduate school unsure of the research they wanted to pursue and needing to find a group during their first year of graduate study. However, as detailed in our previous work [32], this opened the door to a number of new challenges as Carmen searched for a group while trying to adjust to their new environment and taking graduate-level courses; we return to these aspects of Carmen’s story in more detail in Section IV.5.
@@
Our data also revealed how access to advanced coursework and electives was another factor that shaped students’ interests, helping them narrow their research interests sooner. Although Desiree’s interest in high energy theory was confirmed during her research experience, it initially stemmed from taking advanced electives: “During my junior and senior years, I started taking more electives… particle physics, cosmology, things that are not in like the main courseload… [topics] you literally never hear about in undergrad, unless you decide to go take a class on it.” Similarly, Matias cited the graduate-level condensed matter course he took as a senior as being a pivotal influence, as he strongly disliked the undergraduate version. He recalled, “That professor was teaching graduate solid state, and I was like oh shit, here we go again. But I enrolled. And then when we looked at it from a graduate perspective, I started to like it. I was like, wait, okay, actually this is kind of cool. We looked into topics in a deeper sense you know?” Yet students from smaller departments, particularly those without a graduate program, might not offer such courses at all. Collectively, this section indicates some of the ways that a student’s background and access to advanced coursework and undergraduate research opportunities might impact their search for a research group down the road as .
@@ IV.4 Early connections with prospective advisors help students join groups sooner
Figure 5: A summary of codes used in Results Section IV.4. They illustrate three broad stages students described going through while trying to find a research group with congruent research interests. Students generally described a narrowing of choices beginning with entire programs at the school level. Once a few schools were identified, students could more carefully evaluate their alignment with individual research groups within those institutions. A subset of students were specifically interested in the exact project that they would work on if they were to join.
Another major factor that Lovitts cited as predicting success in the PhD was students’ proactivity in searching for a research group [12]. Specifically, she found that in the period before arriving at graduate school, students who collected information about individual faculty advisors were more likely to persist. Students who contacted faculty directly or went for on-campus visits were more likely to complete a PhD than students who relied more on the school’s reputation or looking at faculty websites: “the closer the source of information to the most interpersonal interactional level (faculty), the larger and more significant the difference between completers and noncompleters becomes.” In this section, we illustrate the variation in when students start earnestly looking into research groups, and show how students who engage in the search process earlier tend to join a group more quickly.
@@
Students with broader interests than Wendy indicated that their criteria for narrowing down schools was not as stringent. Isaac said that while applying to schools, “I knew vaguely what area of physics I was interested in… I guess, at the time, particle physics, and maybe like quantum field theory, or something. So my thought was just, you know what, let’s see if there’s anyone out there doing that.” Online resources also played a big role at this time for Isaac, who said he would “go to [the physics department’s] page, look at the professors, and just click on people who seemed interesting.” A common refrain from students at this point in their process was that they wanted to identify programs with several groups whose research seemed to align with their interests. For instance, Blake said that they were drawn to their graduate program because “because there’s a lot of research in what I wanted to do here,” which was appealing because “I didn’t have a super specific focus I wanted within optics.”
@@
Many students ($N=28$) had excerpts coded in the School level alignment category, and students were equally distributed over time to join a group (72% who joined 1st semester or before and 68% who joined after first semester). This is indicative of how regardless of whether students were sure of what research they wanted to pursue, they still went through this process. However, this means 12 students did not discuss looking at their school’s research at all. While several students may have simply omitted this from their narratives, some students specifically noted that they did not filter schools based on research. For instance, two students attended the same graduate school where they went to undergrad, and never contemplated going elsewhere. Another student chose which graduate schools to apply for based on whether they offered a fee waiver, figuring they would sort out their research interests once they arrived. Other students primarily filtered graduate schools based on their location and proximity to family. Meanwhile, several students strongly coupled their choice of graduate school to an individual advisor, and therefore never talked about considering the research going on in the broader department.
@@
Indivdual group level alignment Excerpts with student attempts to determine research interest congruence with just one or a few groups were coded as Evaluating individual group level alignment. While the school level alignment phase commonly had students looking at school websites for help, this phase was most commonly characterized by talking to prospective advisors and graduate students directly. In both stages, students reflected often on their undergraduate research experiences for help deciding whether a research topic seemed interesting. In sum, excerpts from $N=34$ unique students were coded under Evaluating individual group alignment, again indicating that students went through this process regardless of interest level.
@@
Students often described becoming more focused on particular groups once they arrived in graduate school. For instance, Valerie recalled beginning to “go between research group to research group, all these different meetings… definitely comparing them against each other, and then comparing them against the good internship experience I had.” Hassan said that it was “first semester when I slowly started getting more into, like the in-depth portion of the research that [this faculty member] did.” This involved meeting directly with his prospective advisor: “I went and met with him. We briefly talked about what he does, I got a tour of the lab.” Jiya emphasized that one of the primary benefits of being on campus with prospective advisors was that she could ask them directly about their research, saying “I want to hear it from them, like how they describe it.”
@@
Within the category of looking at individual groups, several students described specifically wanting to evaluate research at the Project level. Excerpts coded under this category show how students sought to understand more precisely what they would work on in the lab if they were to join, and what to expect in their day-to-day tasks. For example, after being accepted to his top choice graduate school in his Senior year of undergrad, Alex recalled asking professors, “Where do you see your research going in the next three to six years? Is the work that you’re doing now going to continue, or am I going to join your group and find that you’re pivoting 180 degrees? And you know, now you’re not doing the work that I joined your group to do.” Matias similarly wanted to know what the “day-to-day” of his work would be like in a prospective group, which he was able to answer by earning a fellowship to work there the summer before graduate school. He said, “I would say if I didn’t have that opportunity, I would still be very uncertain right now.” Comparatively fewer students ($N=15$) made statements coded under this sub-category, suggesting that it may not be a topic that students looking for groups consider as often or feel that they are able to evaluate before deciding to join a group.
@@
In total, $N=24$ students talked about evaluating their interest in individual groups before graduate school, while $N=11$ exclusively talked about looking at research groups after arriving in graduate school. Figure 6 plots these two subsets of students, split by when they joined a research group. We see that students who joined a research group during or before their 1st semester were more likely to have evaluated individual groups or projects before they arrived in graduate school than students who found a group second semester or later. 72% ($N=13$) of students who committed to a research group in their first semester or earlier had started the process of evaluating individual research groups before graduate school, compared to 45% ($N=10$) of students who joined second semester or later. Only 11% ($N=2$) of the earlier joiners indicated that they had not already been looking at specific groups prior to arriving in graduate school. On the other hand, 41% ($N=9$) of those who joined second semester or later did not begin closely looking into individual groups until starting their programs. This illustrates the variation in when students start earnestly looking into research groups, and shows how students who engage in the search process earlier tend to join a group more quickly.
Figure 6 also highlights the subset of students in our sample who switched research groups, shown in dark green. The two students represented by the dark green bar on the left are Benjamin and Tabitha, both of whom were also categorized as knowing the exact research they wanted to pursue. Benjamin was guided toward his research group by his undergraduate advisor, who was a collaborator with his prospective graduate advisor; Tabitha’s prospective advisor contacted her directly to recruit her to his group. However, both students found that their graduate labs did not fit them well and switched.
@@ IV.4.1 Connections from undergraduate experiences help students search earlier
The resources that students used while looking at schools or contacting individual research groups varied. Websites helped to figure out which schools might have interesting research, but students often had difficulty gleaning more specific details about a professor’s research solely from online resources. As noted by Irene, websites were “the first point of information. Like when I was first doing research for schools, I would go to the school web page, go to the faculty and look at all their websites. And a lot of them if you go down to the bottom, it says last updated 2015. At that point, you can take none of the information.” For more detailed information, other sources of help like as one-on-one conversations with faculty were significantly more valuable.
@@
At the same time, websites are significantly more accessible and less intimidating than meetings with graduate faculty. Prior to arriving in graduate school, setting up meetings with prospective advisors with whom students are not co-located was more difficult and required more effort. Matias discussed the lack of time associated with finishing undergraduate coursework at the same time as applying to graduate school: “I barely had time because I’m finishing up my semester, I did a double major so every semester was like 20 units, so I was like, up to here with work. And so it was really, it’s really hard to set time aside to look into these things, you know? Also… I was like, I’m not even gonna get in. So what’s the point of looking into it?” Some discussed anxiety associated with reaching out to faculty. For example, Jiya indicated that she struggled with being able to “put myself out there” and start talking to faculty. Isaac recalled that “I guess the biggest question I had was, if I were to summarize, how to find someone I’m interested in and then how to talk to them and have them not think I’m an idiot.”
@@
On the other hand, some students cited their ease of access to faculty at their prospective graduate schools as helpful in navigating these difficulties. For instance, Felipe’s undergraduate research advisor helped him make connections with his eventual group in graduate school: “One of [my mentors] encouraged me to apply here. Because he knew the people in the research group I was interested in working with, so it was like, seems like a nice fit.” Felipe leveraged his undergraduate advisor’s connections to break the ice with his eventual graduate advisor. Similarly, Alex said that emailing faculty was easy after his REU, which he described as “a program whose sole purpose was to help students get into [the school, so I] had those connections essentially made for me. So when I wanted to take advantage, all I needed to do was send an email, basically no introduction necessary.” The networking opportunities that undergraduate research provided allowed some students to more efficiently get in touch with individual groups. Constastingly, Selena cited her background at a small school as a distinct disadvantage during her search. “My school was small enough, we didn’t have connections to these places I was applying. So it wasn’t like I could say to the person I’m working with in my school, can you reach out to someone there? Can you help me find someone? So it was very much like Google, and a lot of clicking, was how I tried to piece it together.”
@@
These results suggest some of the resources used to more closely evaluate research groups (e.g., meetings with faculty) may not be as easily accessible to some students as others. Some are less likely to have the time or connections to look beyond the school level while narrowing down groups, Or, as detailed in our previous work, they simply may not know whether activities such as reaching out to faculty before starting their graduate program are appropriate [50, 32].